If sigma is unknown, X is normal and n=5, then the sampling distribution of standardized X-bar is

  1. t (n-1)
  2. approximately Z
  3. unknown
  4. binomial
Answer #1

solution

If sigma is unknown

the we use t distribution  

t (n-1) =t /2,n-1 = t /2,df

sampling distribution of standardized X-bar is

  1. t (n-1)